OpenStudy (anonymous):

If n+5(n-1)=7,what does n=?

OpenStudy (anonymous):

n=2

jimthompson5910 (jim_thompson5910):

\[\large n+5(n-1)=7\] \[\large n+5n-5=7\] \[\large 6n-5=7\] \[\large 6n=7+5\] \[\large 6n=12\] \[\large n=\frac{12}{6}\] \[\large n=2\]